OpenStudy (anonymous):

How do I do this? Which statement describes the linear function graphed below? A. As x increases by 1, y decreases by 4. B. As x increases by 4, y decreases by 1. C. As x increases by 1, y increases by 4. D. As x increases by 4, y increases by 1.

OpenStudy (anonymous):

Don't give me the answer! I just want someone to teach it to me.

OpenStudy (anonymous):

@hartnn @zepdrix @mathslover @iambatman @blues

OpenStudy (anonymous):

@Kainui @tHe_FiZiCx99 @Luigi0210 @Lena772 @severus4ever

OpenStudy (lena772):

stop mass tagging please.

OpenStudy (marissalovescats):

Well what is the slope? Rise over run. The slope is 1/4 So basically what you're looking at: You move to the right 4 on the x axis (horizontal axis) and you land on 1 on the y axis (the vertical axis) So that tells you: As x increases by 4, y increases by 1.

OpenStudy (anonymous):

@Lena772 I'm not mass tagging. I only tagged ten people.

OpenStudy (marissalovescats):

Another way to see that is At x=4 y is 1 at x=8 y is 2 So you can see when x increases by 4, y increases by 1.

OpenStudy (anonymous):

I see. :) I have a similar one but it's a different kinda graph. Can you help explain it?

OpenStudy (anonymous):

Which statement describes the table below? A. As x increases by 1, y increases by 4. B. As x increases by 1, y increases by 6. C. As x increases by 1, y decreases by 4. D. As x increases by 1, y decreases by 6.

OpenStudy (anonymous):

@marissalovescats

OpenStudy (marissalovescats):

Well whats the difference of: 0 to 1 to 2 to 3? What are you doing to the terms? Whats the difference of: 4 to 10 to 16 ? What are you doing to the terms?

OpenStudy (anonymous):

Your just adding 1. 1+2+3+4+5... Adding 6. 4+6=10+6=16...

OpenStudy (anonymous):

B? @marissalovescats

OpenStudy (marissalovescats):

Yes B :)

OpenStudy (anonymous):

Yay! Thanks!!! :)

OpenStudy (marissalovescats):

You're welcome!

OpenStudy (the_fizicx99):

Muh Bestie @Lena772 is great at math <3

OpenStudy (lena772):

k